# Emergence of Meron Kekulé lattices in twisted Néel antiferromagnets

### v1

@@ Abstract
A Kekulé lattice is an exotic, distorted lattice structure distinguished by alternating bond lengths in contrast to naturally formed atomic crystals. While this structure has been explored through atomic crystals and metamaterials, the possibility of forming a Kekulé lattice from topological solitons in magnetic systems has remained elusive. Here, we propose twisted bilayer easy-plane Néel antiferromagnets as a promising platform for achieving a “Meron Kekulé lattice”—a distorted topological soliton lattice comprised of antiferromagnetic merons as its lattice elements. Using atomistic spin simulations on these magnets, we demonstrate that due to the moiré-induced intricate pattern of interlayer exchange coupling, the cores of these merons are stabilized into the Kekulé-O pattern with different intracell and intercell bond lengths across moiré supercells, hence forming a Meron Kekulé lattice. Furthermore, we showcase that the two bond lengths of the Meron Kekulé lattice can be fine-tuned by adjusting the twist angle and specifics of the interlayer exchange coupling, suggesting extensive control over the meron lattice configuration in contrast to conventional magnetic systems. These discoveries pave the way for exploring topological solitons with distinctive Kekulé attributes, offering intriguing opportunities at the intersection of topological solitons and Kekulé physics.
@@
Merons are particle-like localized spin configurations typically found in easy-plane magnets, which exhibit both out-of-plane polarization in their core regions and in-plane swirling textures away from the cores 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14. These unique characteristics endow merons with distinctive half-integer skyrmion numbers, classifying them as topological solitons 11, 12. The topological stability implied by the skyrmion numbers suggests merons as promising candidates for topological magnetic bits in spintronics, capable of storing and transmitting information 15, 16, 17, 18, 19. However, in continuous magnetic systems, merons typically exist alongside their counterparts, namely antimerons, and often exhibit instability against pair annihilation 6, 7, 14. Consequently, merons may manifest in limited forms such as transient states 6, 7, 14, aggregated groups 5, 8, 9, 10, or a square lattice form as observed in chiral magnet Co8Zn9Mn3 11. The potential instabilities and constrained manifestations of merons hinder the comprehensive study of their fundamental properties and the exploration for future technological applications.
@@
A Kekulé lattice is a distorted lattice structure distinguished by alternating bond lengths. This lattice distortion, often referred to as Kekulé distortion, has been suggested to give rise to a range of intriguing physical phenomena not observed in pristine, undistorted lattices 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 29, 39, 40, 41, 42, 43, 24, 21, 20, 44. Notable examples include charge fractionalization 20, 21, chiral-symmetry breaking 22, 24, 25, 26, 23, 27, 28, and emergence of flat bands 31 in engineered graphene. Additionally, the appearance of topological band structures has been demonstrated in photonic 40, 41 and phononic crystals 42, 43. Experimentally, the Kekulé lattice has been realized in atomic crystals such as graphene through the addition of adatoms 28, 44 and intercalation 30. However, despite the predominant focus on such atomic crystals and metamaterials, the concept of a Kekulé lattice can extend beyond these systems, holding promise for broader applications across various physical systems with distinct lattice elements. Particularly fascinating is the potential creation of a Kekulé lattice from topological solitons in magnetic systems, which would introduce a novel perspective, contrasting with the traditionally observed Bravais lattice forms of topological solitons, such as the triangular lattice seen in the Abrikosov vortex lattice in superconductors 45 and the skyrmion lattice in magnetic systems 46. This intriguing possibility has remained unexplored, representing a pivotal piece yet to be uncovered within the realms of Kekulé physics.
@@
In this study, we discover the creation of a Kekulé lattice from antiferromagnetic merons in twisted bilayer easy-plane Néel antiferromagnets. By conducting atomistic spin simulations on these magnets, we demonstrate that the cores of the antiferromagnetic merons are stabilized due to the moiré-induced spatial modulation of interlayer exchange coupling 47. Notably, the cores of these merons form a honeycomb-lattice-like structure across moiré supercells, exhibiting different intracell and intercell bond lengths (Figure 1). This distorted structure deviates from commonly observed, elementary Bravais-lattice forms 46, 48, 11, 49, 50, yet aligning with the known characteristics of the Kekulé-O pattern 28. Hence, we term our distorted meron lattice a “Meron Kekulé lattice.” We illustrate that the two bond lengths of the Meron Kekulé lattice can be fine-tuned by manipulating the twist angle and specifics of interlayer exchange coupling, offering a high degree of control over the stable meron lattice configuration. These discoveries present a Meron Kekulé lattice as a new classification within Kekulé lattices, introducing a novel pathway to attain this structure in atomic crystals. Moreover, the stability and adjustability of this distorted structure present intriguing possibilities for exploring topological solitons with precise control—a capability rarely achievable within conventional magnetic systems 46, 48, 11, 49, 50, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14.
Results
@@
AFM domain array. We construct twisted bilayer Néel antiferromagnets by rotating two magnetic layers in a honeycomb lattice with a relative twist angle (Figure 2a). These twisted magnets can be described using a Heisenberg spin model given by
$H=$ $\;J\sum_{l=t,b}\sum_{\langle i,j\rangle}\bm{S}_{i}^{l}\cdot\bm{S}_{j}^{l}+A\sum_{l=t,b}\sum_{i}\big{(}\bm{S}_{i}^{l}\cdot\hat{\bm{z}}\big{)}^{2}$
$+\sum_{i,j}J_{ij}^{\perp}\bm{S}_{i}^{t}\cdot\bm{S}_{j}^{b}.$ (1)
Here, $\bm{S}_{i}^{l}$ represents the spin at site $i$ on the top layer ($l=t$) and the bottom layer ($l=b$). The parameter $J=1.0$ meV represents the intralayer AFM exchange interactions between nearest-neighbor spins. $A=0.2$ meV represents the single-ion anisotropy energy favoring in-plane spin orientations. $J_{ij}^{\perp}$ represents the interlayer AFM exchange interactions that modulate as a function of the coordinate displacements $\bm{r}_{ij}^{tb}=\bm{r}_{i}^{t}-\bm{r}_{j}^{b}$ between two spins at $i$- and $j$- sites on the top and bottom layers, respectively. To describe the decaying behavior of $J_{ij}^{\perp}$ as a function of $|\bm{r}_{ij}^{tb}|$, we employ the following exponential function 51, 52, 53, 54, 55:
@@
Here, the parameter $J_{0}^{\perp}$ represents the maximum value of $J_{ij}^{\perp}$ at $|\bm{r}_{ij}^{tb}|=d$, where $d$ denotes the perpendicular interlayer distance; the parameter $\alpha$ describes the decay rate of $J_{ij}^{\perp}$ with respect to the increase of $|\bm{r}_{ij}^{tb}|$. We choose $\alpha=15$ with the consideration that $J_{ij}^{\perp}$ decreases to approximately 10% of $J_{0}^{\perp}$ at the second-nearest neighbor distance between the top and bottom layers in the AA stacking (i.e., $|\bm{r}_{ij}^{tb}|=\sqrt{d^{2}+a^{2}}$), where $a$ is the nearest-neighbor distance in a honeycomb lattice. In principle, one can calculate the value of $\alpha$ through ab initio calculations 51. Additionally, we take $J^{\perp}_{0}=0.3~{}\textrm{meV}$, $d=7~{}\textrm{\AA{}}$, and $a=4~{}\textrm{\AA{}}$, aligned with typical values observed in diverse vdW magnetic materials 56. We use the twist angle $\theta=1.61$°, which results in a moiré supercell with 5044 sites. These parameter values are utilized uniformly throughout this study unless explicitly specified.
The moiré superlattice of the twisted bilayer encompasses various local stacking patterns, such as AA, AB, and BA, within its supercell (Figure 2a). This stacking modulation leads to diverse spin alignments between the top and bottom layers. Two characteristic alignment patterns emerge:
@@
In the AA stacking regions (red patches in Figure 2a), the A-type order predominates since spins on the same sublattice in the top and bottom layers (e.g., $\bm{S}_{i,A}^{t}$ and $\bm{S}_{i,A}^{b}$ for the sublattice A) are coupled through the AFM coupling $J_{ij}^{\perp}$, as illustrated in the left panel of Figure 2b. Conversely, the AB stacking regions (blue patches in Figure 2a) favor the B-type order, as $\bm{S}_{i,A}^{t}$ and $\bm{S}_{i,B}^{b}$ are primarily coupled, while $\bm{S}_{i,B}^{t}$ and $\bm{S}_{i,A}^{b}$ remain effectively decoupled due to their large separations, as depicted in the middle panel of Figure 2b. Similarly, the BA stacking regions (cyan patches in Figure 2a) prefer a parallel alignment with similar considerations, as shown in the right panel of Figure 2b. We illustrate the intricate coexistence of these two orders using a map of the local interlayer exchange energy $J_{i}^{\perp}=\sum_{j}J_{ij}^{\perp}\bm{S}_{i}^{t}\cdot\bm{S}_{j}^{b}$, computed for the B-type Néel order where $\bm{S}_{i,A}^{t}=\bm{S}_{i,A}^{b}=(1,0,0)$ and $\bm{S}_{i,B}^{t}=\bm{S}_{i,B}^{b}=(-1,0,0)$. The resulting map, presented in Figure 2c, showcases three distinct local regions: (i) “AFM patches” preferring the A-type order, as indicated by their AFM coupling character between two Néel vectors (i.e., $J_{i}^{\perp}>0$), (ii) “FM patches” preferring the B-type order, as indicated by their FM coupling character (i.e., $J_{i}^{\perp}<0$), and (iii) a neutral intermediate region lacking any specific preferred order, exhibiting negligible coupling (i.e., $J_{i}^{\perp}\approx 0$).
@@
Our atomistic spin simulations on Eq. (Emergence of Meron Kekulé lattices in twisted Néel antiferromagnets), conducted using an iterative optimization method 55, 54, 47, reveal a zero temperature magnetic phase diagram shown in Figure 2d (for detailed methods, refer to the Methods). This diagram displays two distinct magnetic phases: AFM and magnetic domain (MD) phases. The AFM phase displays a uniform spin configuration characterized by the B-type Néel order. In contrast, the MD phase exhibits a nonuniform configuration (Figure 2eg), where the A-type Néel order emerges within each AFM patch, forming circular-shaped domains, whereas the B-type Néel order persists outside these patches (Figure 2g). We term this spin configuration an “AFM domain array,” 47 attributed to the AFM-order-like configuration (i.e., antiparallel alignment of the Néel vectors between the two layers) within these domains that are arranged in an array pattern.
The AFM domain array configuration minimizes $J_{i}^{\perp}$ across both the AFM and FM patch regions. This contrasts with the uniform AFM configuration, which only achieves low energy within the FM patches, leaving the AFM patches with higher energy. Our finding reveals that this energy reduction outweighs an increase from the domain wall energy, leading to an energy-minimized ground state configuration (for detailed information, refer to the Supplementary Information). A previous effective continuum model calculation suggests that the lowered interlayer exchange energy and domain wall energy are given by $\Delta E_{\textrm{inter}}\sim-\bar{J}_{\perp}\theta^{-2}$ and $\Delta E_{\textrm{intra}}\sim+J$, respectively, where $\bar{J}_{\perp}$ denotes the average value of $J_{i}^{\perp}$ within an AFM patch 54, 55, 47. Notably, $|\Delta E_{\textrm{inter}}|$ outweighs $|\Delta E_{\textrm{intra}}|$ in a small twist angle regime ($\theta\leq\theta_{c}\sim\sqrt{\bar{J}_{\perp}/J}$). Furthermore, the formula $\theta_{c}=C\sqrt{\bar{J}_{\perp}/J}$ fits well the observed phase boundary with the parameters $\bar{J}_{\perp}=0.3$ meV and $C=23.24$ (the dashed line in Figure 2d). Consequently, the AFM domain array of the MD phase becomes a ground state in this small twist angle regime.
@@
Meron Kekulé lattice. We demonstrate that the AFM domain array, depicted in Figure 2eg, can accommodate merons as topological defects around its domain boundary, leading to a Meron Kekulé lattice. Figure 3 presents a stable spin configuration obtained through the relaxation of a random initial configuration. On the top layer, the Néel vectors exhibit six-fold in-plane swirling textures within a supercell, which are labeled by $\textrm{M}_{1}$, $\textrm{M}_{2}$, $\textrm{M}_{3}$, $\overline{\textrm{M}}_{1}$, $\overline{\textrm{M}}_{2}$, and $\overline{\textrm{M}}_{3}$ (Figure 3a). The vorticity density of these textures is computed as 57:
$v(x,y)=\frac{1}{\pi}\hat{\bm{z}}\cdot(\partial_{x}\bm{N}(x,y)\times\partial_{y}\bm{N}(x,y)),$ (5)
@@
The six vortices exhibit out-of-plane polarization in their core regions (Figure 3a), indicating that they carry an additional topological number defined in three-dimensional spin space, known as a skyrmion number. The local density profile for the skyrmion number is calculated as 11, 12:
$q(x,y)=\frac{1}{4\pi}\bm{N}(x,y)\cdot(\partial_{x}\bm{N}(x,y)\times\partial_{y}\bm{N}(x,y)).$ (6)
[…]
The cores of the merons and antimerons form six vertices of a hexagon along the boundary of the AFM patch, with each pair, ($\textrm{M}_{1}$, $\overline{\textrm{M}}_{1}$), ($\textrm{M}_{2}$, $\overline{\textrm{M}}_{2}$), and ($\textrm{M}_{3}$, $\overline{\textrm{M}}_{3}$), positioned oppositely among these vertices (Figure 3a). In this arrangement, the merons cancel out their swirling textures away from the cores, maintaining a localized configuration within a supercell. We refer to this distinctive local configuration as a “Meron hexad.” Despite its intricate texture, the Meron hexad attains an AFM domain array configuration that minimizes interlayer exchange energy across the bulk region, including both FM and AFM patches (Figure 3e). This is achieved through two key mechanisms: (i) a 180° rotation of the Néel vectors across FM and AFM patches in the top layer facilitated by the hexagonal alignment of the meron cores around the AFM patch (Figure 3a) and (ii) a nearly uniform alignment of Néel vectors with a slight tilt in opposing directions between FM patches in the bottom layer (Figure 3b). Consequently, the Meron hexad attains a minimized energy (e.g., $-1.5176$ meV per spin for the configuration presented in Figure 3), comparable to that of the MD state (e.g., $-1.5128$ meV per spin for the configuration presented in Figure 2ef). Our findings reveal that the small energy gap between the two states primarily stems from the core energy due to intralayer exchange interactions. Further details can be found in the Supplementary Information.
@@
Figure 4: Manipulation of the Kekulé structure through parameter adjustments. a Schematic illustration of the Meron Kekulé lattice structure. Yellow dots denote the position of the meron cores. Black lines denote intracell bonds between meron cores within a Meron hexad, while light gray lines denote intercell bonds between meron cores across different Meron hexads. $l$ and $l^{\prime}=l-2L$ denote the intracell and intercell bond lengths, respectively. $r$ denotes the size of AFM patches marked by puple circles. bd Manipulation of the Kekulé lattice structure through three different parameters: b the moiré superlattice size ($L$), c the decay rate parameter of interlayer exchange interactions $J_{ij}^{\perp}$ ($\alpha$), and d the maximum strength of $J_{ij}^{\perp}$ ($J_{0}^{\perp}$). In b, the inset shows the value of the Kekulé constant ($\lambda=\frac{l^{\prime}-l}{L}$). The dashed lines represent empirical fitting curves. In c, the insets display $J_{i}^{\perp}$ for $\alpha=10$ (left) and $\alpha=30$ (right), respectively. The same color scale used in Figure 2c is applied in the insets. In d, the dashed line represents the fitting curve for the $l$ values. e Effective energy function $E(l)$ for the MK state, plotted in a normalized form: $F(l/L)=\frac{E(l)-\textrm{min}[E(l)]}{\textrm{max}[E(l)]-\textrm{min}[E(l)]}$. Here, ‘min’ and ‘max’ indicate the minimum and maximum values, respectively. Different colors indicate distinct values of $\log_{10}(p)$, where $p$ represents the dimensionless ratio between the Coulomb interactions and the moiré field, which represent the two effective interactions influencing merons in the effective model. The parameter $r/L=0.25$ is used. f Illustration of the stabilization condition of the MK state as a function of $L$ and $J_{0}^{\perp}$. The MK state is stable in the blue region while unstable in the devoid region. The markers represent simulation data points, while the dashed line is a fitting curve defined as $J_{0}^{\perp}=40/L^{2}$.

### v2

@@ Abstract
A Kekulé lattice is an exotic, distorted lattice structure exhibiting alternating bond lengths, distinguished from naturally formed atomic crystals. Despite its evident applicability, the formation of a Kekulé lattice from topological solitons in magnetic systems has remained elusive. Here, we propose twisted bilayer easy-plane Néel antiferromagnets as a promising platform for achieving a “Meron Kekulé lattice”—a distorted topological soliton lattice comprised of antiferromagnetic merons as its lattice elements. We demonstrate that the cores of these merons are stabilized into the Kekulé-O pattern with different intracell and intercell bond lengths across moiré supercells, thereby forming a Meron Kekulé lattice. Moreover, the two bond lengths of the Meron Kekulé lattice can be fine-tuned by adjusting the twist angle and specifics of the interlayer exchange coupling, suggesting extensive control over the meron lattice configuration in contrast to conventional magnetic systems. These discoveries pave the way for exploring topological solitons with distinctive Kekulé attributes.
@@
A Kekulé lattice is a distorted lattice structure distinguished by alternating bond lengths. This lattice distortion has been suggested to give rise to intriguing physical phenomena not observed in undistorted lattices 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 10, 20, 21, 22, 23, 24, 5, 2, 1, 25. Notable examples include charge fractionalization 1, 2, chiral-symmetry breaking 3, 5, 6, 7, 4, 8, 9, and emergence of flat bands 12 in graphene. Additionally, the appearance of topological band structures has been demonstrated in photonic 21, 22 and phononic crystals 23, 24. Experimentally, the Kekulé lattice has been realized in graphene 9, 25, 11. However, the concept of a Kekulé lattice can extend beyond these systems, holding promise for broader applications across various physical systems with distinct lattice elements. Particularly fascinating is the potential creation of a Kekulé lattice from topological solitons in magnetic systems, which would introduce a novel perspective, contrasting with the traditionally observed Bravais lattice forms of topological solitons, such as the triangular lattice seen in the Abrikosov vortex lattice in superconductors 26 and the skyrmion lattice in magnetic systems 27. This intriguing possibility has remained unexplored, representing a pivotal piece yet to be uncovered within the realms of Kekulé physics.
@@
In this study, we discover the creation of a Kekulé lattice from antiferromagnetic merons in twisted bilayer easy-plane Néel antiferromagnets. By conducting atomistic spin simulations on these magnets, we demonstrate that the cores of the antiferromagnetic merons are stabilized due to the moiré-induced spatial modulation of interlayer exchange coupling 28. Notably, the cores of these merons form a honeycomb-lattice-like structure across moiré supercells, exhibiting different intracell and intercell bond lengths (Figure 1). This distorted structure deviates from commonly observed Bravais-lattice forms 27, 29, 30, 31, 32, yet aligning with the known characteristics of the Kekulé-O pattern 9. Hence, we term our distorted meron lattice a “Meron Kekulé lattice.” We illustrate that the two bond lengths of the Meron Kekulé lattice can be fine-tuned by manipulating the twist angle and specifics of interlayer exchange coupling, offering a high degree of control over the stable meron lattice configuration. These discoveries present a Meron Kekulé lattice as a new classification within Kekulé lattices, introducing a novel pathway to attain this structure. Moreover, the stability and adjustability of this distorted structure present intriguing possibilities for exploring topological solitons with precise control—a capability rarely achievable within conventional magnetic systems 27, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 30, 43, 44, 45.
@@
We construct twisted bilayer Néel antiferromagnets by rotating two magnetic layers in a honeycomb lattice with a relative twist angle (Figure 2a). These twisted magnets can be described using a Heisenberg spin model given by
$H=$ $\;J\sum_{l=t,b}\sum_{\langle i,j\rangle}\bm{S}_{i}^{l}\cdot\bm{S}_{j}^{l}+A\sum_{l=t,b}\sum_{i}\big{(}\bm{S}_{i}^{l}\cdot\hat{\bm{z}}\big{)}^{2}$ (1)
$+\sum_{i,j}J_{ij}^{\perp}\bm{S}_{i}^{t}\cdot\bm{S}_{j}^{b}.$
Here, $\bm{S}_{i}^{l}$ represents the spin at site $i$ on the top layer ($l=t$) and the bottom layer ($l=b$). The parameter $J=1.0$ meV represents the intralayer AFM exchange interactions between nearest-neighbor spins. $A=0.2$ meV represents the single-ion anisotropy energy favoring in-plane spin orientations. $J_{ij}^{\perp}$ represents the interlayer AFM exchange interactions that modulate as a function of the coordinate displacements $\bm{r}_{ij}^{tb}=\bm{r}_{i}^{t}-\bm{r}_{j}^{b}$ between two spins at $i$- and $j$- sites on the top and bottom layers, respectively. To describe the decaying behavior of $J_{ij}^{\perp}$ as a function of $|\bm{r}_{ij}^{tb}|$, we employ the following exponential function 46, 47, 48, 49, 50:
@@
Here, the parameter $J_{0}^{\perp}$ represents the maximum value of $J_{ij}^{\perp}$ at $|\bm{r}_{ij}^{tb}|=d$, where $d$ denotes the perpendicular interlayer distance; the parameter $\alpha$ describes the decay rate of $J_{ij}^{\perp}$ with respect to the increase of $|\bm{r}_{ij}^{tb}|$. The twist angle $\theta$ defines the size of the moiré superlattice $L$ through the relationship: $L\approx\frac{\sqrt{3}a}{\theta}$ ($\theta$: in radians). We choose $\alpha=15$, $J^{\perp}_{0}=0.3~{}\textrm{meV}$, $d=7~{}\textrm{\AA{}}$, and $a=4~{}\textrm{\AA{}}$, aligned with typical values observed in diverse vdW magnetic materials 51. We use the twist angle $\theta=1.61$°. These parameter values are utilized throughout this study unless explicitly specified.
The moiré superlattice of the twisted bilayer encompasses various local stacking patterns, such as AA, AB, and BA, within its supercell (Figure 2a). This stacking modulation leads to diverse spin alignments between the top and bottom layers 46. Two characteristic alignment patterns emerge:
@@
In the AA stacking regions, the A-type order predominates since spins on the same sublattice in the top and bottom layers (e.g., $\bm{S}_{i,A}^{t}$ and $\bm{S}_{i,A}^{b}$ for the sublattice A) are coupled through the AFM coupling $J_{ij}^{\perp}$ (the left panel of Figure 2b). Conversely, the AB stacking regions favor the B-type order, as $\bm{S}_{i,A}^{t}$ and $\bm{S}_{i,B}^{b}$ are primarily coupled, while $\bm{S}_{i,B}^{t}$ and $\bm{S}_{i,A}^{b}$ remain effectively decoupled due to their large separations (the middle panel). The BA stacking regions prefer the B-type order with similar considerations (the right panel). The map of local interlayer exchange energy, computed as $J_{i}^{\perp}=\sum_{j}J_{ij}^{\perp}\bm{S}_{i}^{t}\cdot\bm{S}_{j}^{b}$, where $\bm{S}_{i,A}^{t}=\bm{S}_{i,A}^{b}=(1,0,0)$ and $\bm{S}_{i,B}^{t}=\bm{S}_{i,B}^{b}=(-1,0,0)$ reveals three distinct local regions (Figure 2c): (i) “AFM patches” preferring the A-type order, as indicated by their AFM coupling character between two Néel vectors (i.e., $J_{i}^{\perp}>0$), (ii) “FM patches” preferring the B-type order, as indicated by their FM coupling character (i.e., $J_{i}^{\perp}<0$), and (iii) a neutral intermediate region lacking any specific preferred order, exhibiting negligible coupling (i.e., $J_{i}^{\perp}\approx 0$).
@@
Our atomistic spin simulations on Eq. (1), conducted using an iterative optimization method 50, 49, 28, reveal a zero temperature magnetic phase diagram shown in Figure 2d. For detailed methods, refer to Supporting Information (SI) 52. This diagram displays two distinct magnetic phases: AFM and magnetic domain (MD) phases. The AFM phase displays a uniform spin configuration characterized by the B-type Néel order. In contrast, the MD phase exhibits a nonuniform configuration (Figure 2eg), where the A-type Néel order emerges within each AFM patch, forming circular-shaped domains, whereas the B-type Néel order persists outside these patches (Figure 2g). We term this spin configuration an “AFM domain array” 28. The AFM domain array configuration minimizes $J_{i}^{\perp}$ across both the AFM and FM patch regions. This energy reduction outweighs domain wall energy in a small twist angle regime $\theta<\theta_{c}$, where $\theta_{c}=C\sqrt{\bar{J}_{\perp}/J}$ and $\bar{J}_{\perp}=0.3$ meV and $C=23.24$ 49, 50, 28, resulting in the emergence of the MD phase as the ground state (Figure 2d).
@@
We demonstrate that the AFM domain array, depicted in Figure 2eg, can accommodate merons as topological defects around its domain boundary, leading to a Meron Kekulé lattice. Merons are vortexlike topological solitons carrying half-skyrmion numbers 30, 43, typically found in easy-plane magnets 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 30, 43, 44, 45. The local density profile for the skyrmion number is calculated as:
$q(x,y)=\frac{1}{4\pi}\bm{N}(x,y)\cdot(\partial_{x}\bm{N}(x,y)\times\partial_{y}\bm{N}(x,y)),$ (5)
where $\bm{N}(x,y)$ represents the normalized interpolation of the Néel vectors. Furthermore, in-plane swirling textures away from the cores of merons are characterized by vorticity. The local density profile for the vorticity is calculated as 53:
$v(x,y)=\frac{1}{\pi}\hat{\bm{z}}\cdot(\partial_{x}\bm{N}(x,y)\times\partial_{y}\bm{N}(x,y)).$ (6)
@@
Figure 4: Manipulation of the Kekulé structure through parameter adjustments. a Schematic illustration of the Meron Kekulé lattice structure. Yellow dots denote the position of the meron cores. Black lines denote intracell bonds between meron cores within a Meron hexad, while light gray lines denote intercell bonds between meron cores across different Meron hexads. $l$ and $l^{\prime}=l-2L$ denote the intracell and intercell bond lengths, respectively. $r$ denotes the size of AFM patches marked by puple circles. bd Manipulation of the Kekulé lattice structure through three different parameters: b the moiré superlattice size ($L$), c the decay rate parameter of interlayer exchange interactions $J_{ij}^{\perp}$ ($\alpha$), and d the maximum strength of $J_{ij}^{\perp}$ ($J_{0}^{\perp}$). In b, the inset shows the value of the Kekulé constant ($\lambda=\frac{l^{\prime}-l}{L}$). The dashed lines represent empirical fitting curves, described by $r=0.28L+0.05$, $l=0.31L-1.04$, and $l^{\prime}=0.38L+2.09$. In c, the insets display $J_{i}^{\perp}$ for $\alpha=10$ (left) and $\alpha=30$ (right), respectively. The same color scale used in Figure 2c is applied in the insets. In d, the dashed line represents the fitting curve for the $l$ values. e Effective energy function $E(l)$ for the MK state, plotted in a normalized form: $F(l/L)=\frac{E(l)-\textrm{min}[E(l)]}{\textrm{max}[E(l)]-\textrm{min}[E(l)]}$. Here, ‘min’ and ‘max’ indicate the minimum and maximum values, respectively. Different colors indicate distinct values of $\log_{10}(p)$, where $p$ represents the dimensionless ratio between the Coulomb interactions and the moiré field, which represent the two effective interactions influencing merons in the effective model. The parameter $r/L=0.25$ is used. f Illustration of the stabilization condition of the MK state as a function of $L$ and $J_{0}^{\perp}$. The MK state is stable in the blue region while unstable in the devoid region. The markers represent simulation data points, while the dashed line is a fitting curve defined as $J_{0}^{\perp}=40/L^{2}$.